# Do Text-to-Vis Benchmarks Test Real Use of Visualisations?

### v1

@@ 1 Introduction
In this study, we gathered publicly available code from the Stack [^1] A 6TB collection of open source code from GitHub Kocetkov et al. (2022) to analyse human preferences in making visualisations using libraries across four programming languages: Python, R, Javascript, and Vega. Since each library has different names for the same visualisation types and properties, we extracted key visualisation code and developed a cross-language mapping for several hundred functions and arguments.222 For example, a bar plot is produced with bar() and barh() in Python, but barplot() in R. Our data, code, and mappings are published at https://github.com/giahy2507/text-to-vis-benchmarks-assessment.
Using this aligned data, we analysed to discover user behaviours when making visualisation and identified similarities and differences between real-world and benchmark datasets. Our analysis considered the chart types, functions called to define properties, and the arguments that modify how those functions behave. We observed that existing benchmarks tend to focus on one aspect of the text-to-vis challenge, including code synthesis, data presentation, or aesthetic attribute adjustment. Additionally, while one dataset shows potential for advancing an end-to-end benchmark, the remaining two diverge significantly from real-world scenarios. This underscores the necessity for new benchmarks that not only reflect the trends identified in our study but also encompass all dimensions of the problem. Only then will we be able to measure our progress on this valuable and challenging task.
@@ 3 Data Collection
As for benchmark datasets, we examined three publicly available ones, including nvBench Luo et al. (2021), ChartDialog Shao and Nakashole (2020), and PlotCoder Chen et al. (2021). They vary in settings and scales, as described in Table 1 and 2. While nvBench and ChartDialog serve as end-to-end text-to-vis benchmarks, involving prompts and plotted data as input as well as visualisation as output, PlotCoder is purely a code synthesis dataset, lacking two of these essential components. Appendix B illustrates examples for these datasets.
@@ 4 Cross-language Mapping Table
To compare the data described in previous sections, we established a cross-language mapping table, based on frequently used parameters. This involved selecting the top 500 frequently used parameters, identifying categories and attributes, and finally, clarifying correctness based on the libraries’ documentation and code execution. Ultimately, it took 40 hours to make the table, comprising 8 categories, 62 attributes, and around 850 parameters across 4 visualisation languages. Figure 1 briefly shows an attribute named “x-axis title,” while details can be found in Appendix C.
@@ 6 Conclusion
In this paper, we analyzed whether text-to-vis benchmarks accurately reflect real-world usage by presenting several analyses on chart types, frequent attributes, and program complexity. Our results show that only one of the three benchmarks is significantly aligned with real-world applications. However, this dataset needs significant modification to be an end-to-end benchmark, which receives prompt and plotted data as input and produces visualization as output. This entails additional efforts such as collecting input data, ensuring code executability, and reviewing visual output, as discussed in Section 5.2. Moreover, we provided insights for future benchmark development, suggesting the evaluation of relevant attributes and challenging charts that better reflect end users’ preferences. We believe that if a system can perform successfully on a benchmark that includes all the cases we observed, it will likely perform well in most real-world scenarios.
@@
This study offers analyses of datasets and acknowledges several limitations. Firstly, our examination was restricted to only four visualisation libraries, each corresponding to a different programming language. This narrow scope may not adequately capture the diversity of applications and use cases within the field. Although we attempted to analyse MatLab code files in The Stack dataset, they are miscategorized in The Stack, processed with the wrong extension. [^9] https://huggingface.co/datasets/bigcode/the-stack-dedup/blob/main/programming-languages.json Despite our efforts to clarify this issue by reaching out to the project authors, we have yet to receive a response. Secondly, our investigation is based on public code, mainly representing programmers with different visualisation levels, including novices, practitioners, and experts. If the target users in a visualisation application like Tableau  [^10] https://www.tableau.com/, the level of validity of our research will decrease. Lastly, this study concludes with an analysis and assessment of existing benchmark datasets without proposing solutions. Nevertheless, we believe that the insights and recommendations provided in this work are valuable for any dataset maker and future studies.
@@ Appendix C Cross-language Mapping Table
Due to the length, we only include a part of the table in Table 4. The whole one can be found in our repository at https://github.com/giahy2507/text-to-vis-benchmarks-assessment.

### v3

@@ 1 Introduction
In this study, we gathered publicly available code from the Stack [^1] A 6TB collection of open source code from GitHub Kocetkov et al. (2022) to analyse human preferences in making visualisations using libraries across four programming languages: Python, R, Javascript, and Vega. Since each library has different names for the same visualisation types and properties, we extracted key visualisation code and developed a cross-language mapping for several hundred functions and arguments.222 For example, a bar plot is produced with bar() and barh() in Python, but barplot() in R. Our data and code are at https://github.com/giahy2507/text-to-vis-bench-assessment.
Using this aligned data, we analysed user behaviours when making visualisation and identified similarities and differences between real-world and benchmark datasets. Our analysis considered the chart types, functions called to define properties, and the arguments that modify how those functions behave. We observed that (1) existing benchmarks tend to focus on only one aspect of the text-to-vis challenge, either code synthesis, data presentation, or aesthetic attribute adjustment, and (2) only one of the datasets is consistent with real-world use, and that dataset is limited by its lack of executability of code outputs.
Success on existing benchmarks does not mean systems are useful for real-world use. We need new benchmarks that cover all aspects of the problem and are consistent with the patterns of use. Only then will we be able to measure our progress on this valuable and challenging task.
@@ 3 Data Collection
As for benchmark datasets, we examined three publicly available ones, including nvBench Luo et al. (2021), ChartDialog Shao and Nakashole (2020), and PlotCoder Chen et al. (2021). They vary in settings and scales, as described in Table 1 and 2. While nvBench and ChartDialog serve as end-to-end text-to-vis benchmarks, involving prompts and plotted data as input as well as visualisation as output, PlotCoder is purely a code synthesis dataset, lacking two of these essential components. Appendix B depicts their examples.
@@ 4 Cross-language Mapping Table
To compare the data described in previous sections, we established a cross-language mapping table, based on frequently used parameters. This involved selecting the top 500 frequently used parameters, identifying categories and attributes, and finally, clarifying correctness based on the libraries’ documentation and code execution. Ultimately, the mapping table comprises 8 categories, 62 attributes, and around 850 parameters across 4 visualisation languages. Figure 1 briefly shows an attribute named “x-axis title,” while details can be found in Appendix C.
@@ 6 Conclusion
In this paper, we analyzed whether text-to-vis benchmarks accurately reflect real-world usage by presenting analyses of chart types, frequent attributes, and program complexity. Our results show that only one of the standard three benchmarks is aligned with real-world use. That dataset has its own critical limitation: it cannot be used as an end-to-end benchmark, going from a request and data as input to a visualization as output. As well as critiquing current benchmarks, we provide guidance for future benchmark development, suggesting the evaluation of relevant attributes and challenging charts that better reflect end users’ preferences. Such a benchmark would guide the development of useful and impactful systems.
@@
This study offers analyses of datasets and acknowledges several limitations. Firstly, our examination was restricted to only four visualisation libraries, each corresponding to a different programming language. This narrow scope may not adequately capture the diversity of applications and use cases within the field. Although we attempted to analyse MatLab code files in The Stack dataset, they are miscategorised in The Stack, processed with the wrong extension. [^9] https://huggingface.co/datasets/bigcode/the-stack-dedup/blob/main/programming-languages.json Despite our efforts to clarify this issue by reaching out to the project authors, we have yet to receive a response. Secondly, our investigation is based on public code, mainly representing programmers with different visualisation levels, including novices, practitioners, and experts. If the target users in a visualisation application like Tableau  [^10] https://www.tableau.com/, the level of validity of our research will decrease. Lastly, this study concludes with an analysis and assessment of existing benchmark datasets without proposing solutions. Nevertheless, we believe that the insights and recommendations provided in this work are valuable for any dataset maker and future studies.
@@ Appendix C Cross-language Mapping Table
Due to the length, we only include a part of the table in Table 4. The whole one can be found in our repository at https://github.com/giahy2507/text-to-vis-bench-assessment.